# Three Fundamental Questions in Modern Infinite-Domain Constraint Satisfaction

### v2

@@ 1.1. The finite-domain CSP dichotomy theorem
Let $\mathfrak{A}$ be a finite relational structure. Then either $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$ and $\mathrm{CSP}(\mathfrak{A})$ is NP-complete, or $\mathrm{Pol}(\mathfrak{A})$ satisfies the Siggers identity
@@ 1.2. The infinite-domain CSP tractability conjecture
  • $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$ (and consequently, $\mathrm{CSP}(\mathfrak{A})$ is NP-complete);
@@ 2.1. Algebraicity is irrelevant
    • $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ pp-constructs $\mathfrak{K}_{3}$ if and only if $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$.
@@ 2.3. Everything is a cheese
    $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$ if and only if $\mathfrak{A}^{\prime}$ pp-constructs $\mathfrak{K}_{3}$.
@@
    Every finite cheese for $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ pp-constructs $\mathfrak{K}_{3}$; hence, $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ is not finitely tractable.
@@ 3. Basic definitions
The set $\{1,\dots,n\}$ is denoted by $[n]$, and we use the bar notation $\bar{t}$ for tuples. For a function $f\colon A^{n}\rightarrow B$, the component-wise action of $f$ on $k$-tuples $(x_{1,1},\dots,x_{1,k}),$ $\dots$,… , $(x_{n,1},\dots,x_{n,k})\in A^{k}$ is defined as the $k$-tuple
@@ 4.1. Injectivity and Datalog reductions
Datalog reductions between CSPs are typically specified using factorisation-free Datalog-interpretations with parameters [1, 38], which are a particular case of logical interpretations with parameters. In the present paper, we only need a very basic understanding of Datalog-reductions, which can be easily explained on an informal level. Datalog is defined by adding formation rules to the existential positive fragment of first-order logic whose semantics is defined with inflationary fixed-points of definable operators. Every existential positive first-order formula is a Datalog formula and, if $\phi(\bar{x})$ is a Datalog formula over some relational signature $\tau\cup\{R\}$ with $R\notin\tau$, then $[\mathsf{ifp}_{R}\,\phi](\bar{x})$ is a Datalog formula over the signature $\tau$ whose semantics is given as follows. For a $\tau$-structure $\mathfrak{A}$ and a tuple $\bar{a}$ over $A$ matching the arity of $\bar{x}$, say $k$, we have $\mathfrak{A}\models[\mathsf{ifp}_{R}\,\phi](\bar{a})$ if and only if $\bar{a}$ is contained in the inflationary fixed-point of the operator $F^{\mathfrak{A}}_{\phi,R}(X)\coloneqq\{\bar{x}\in A^{k}\mid\mathfrak{A}_{X}\models\phi(\bar{x})\}$, i.e., the limit of the sequence $X_{0}\coloneqq\emptyset$ and $\smash{X_{i+1}\coloneqq X_{i}\cup F^{\mathfrak{A}}_{\phi,R}(X_{i})}$. For example, the Datalog formula
@@
computes the transitive closure of $<$.
For a relational $\tau$-structure $\mathfrak{A}$, we say that $\mathrm{CSP}(\mathfrak{A})$ is solvable in Datalog if there exists a Datalog sentence defining the class of all finite $\tau$-structures which do not homomorphically map to $\mathfrak{A}$. For example, $\mathrm{CSP}(\mathbb{Q};<)$ is solvable in Datalog using the sentence $\exists u\ldotp\phi_{<}^{tc}(u,u)$. It is not hard to see that every Datalog formula is specified by a finite set of rules of the form $R(\bar{x})\Leftarrow R_{1}(\bar{x}_{1})\wedge\cdots\wedge R_{m}(\bar{x}_{m})$ where $R$ is a fixed-point variable. In the case of the above Datalog formula computing the transitive closure of $<$, for example, the rules are $T(x,z)\Leftarrow(x<z)$ and $T(x,z)\Leftarrow T(x,y)\wedge T(y,z)$. When it comes to Datalog-reductions between CSPs, we only need the following intuitive understanding of this concept.
@@ 4.2. Wreath products
Consider the wreath product $(\mathbb{Q};<_{1})\wr(\mathbb{Q};<_{2})$ of two isomorphic copies of $(\mathbb{Q};<)$ with disjoint signatures. The union $<_{\text{conv}}$ of $<_{1}$ and $<_{2}$ in this structure, together with $E$, is the random convexly ordered equivalence relation from Corollary 6.8 in [52].
@@ 4.3. Removing algebraicity
Let $\tau$ and $\sigma$ be the signatures of $\mathfrak{A}$ and $\mathfrak{B}$, respectively. We first define the auxiliary structure $\mathfrak{B}^{\prime}\coloneqq(\mathbb{Q};<)\wr\mathfrak{B}$. As shown in Examples 1.3 and 2.1, the structure $(\mathbb{Q};<)$ is finitely bounded, homogeneous, and has no algebraicity. It then follows immediately from Proposition 4.5(2) that $\mathfrak{B}^{\prime}$ has no algebraicity. The expansion $\mathfrak{B}^{\prime}_{\neq}$ of $\mathfrak{B}^{\prime}$ by the binary inequality has no algebraicity because taking expansions by first-order definable relations does not change the automorphism group; the same is true for the structure $\mathfrak{B}^{\prime}_{\neq,I_{4}}$. Note that removing all symbols in $\sigma\setminus\tau$ as well as the symbol $<$ from $\mathfrak{B}^{\prime}_{\neq}=\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ yields $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
@@
We claim that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective. To this end, let $\mathfrak{J}$ be an arbitrary finite structure for which there exists a homomorphism $h\colon\mathfrak{J}\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We define $\mathfrak{J}_{\pi}$ as the $(\tau\cup\{E\})$-structure with domain $J$ such that, for every $R\in\tau\cup\{E\}$, we have $\bar{t}\in R^{\mathfrak{J}_{\pi}}$ if and only if $h(\bar{t})\in R^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$. It is easy to see that $\mathfrak{J}_{\pi}$ embeds into the $(\tau\cup\{E\})$-reduct of $\mathfrak{B}^{\prime}$ (note that if multiple elements in $J$ are mapped to $(q,a)$ for some $q\in\mathbb{Q}$ and $a\in A$, we can always map them to different elements in ×$\mathbb{Q}\times\{a\}$). Hence, $\mathfrak{J}_{\pi}$ expanded by the binary inequality predicate embeds into $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and admits an injective homomorphism from $\mathfrak{J}$. We conclude that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective.
@@
Next, we informally describe two Datalog-reductions $\mathcal{I}^{\ast}$ and $\mathcal{I}$: from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$ and back, starting with the former. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$. Denote by $\sim$ the equivalence closure of $E^{\mathfrak{J}}$, and let $q_{\sim}\colon\mathfrak{J}\rightarrow(\mathfrak{J}/{\sim}),x\mapsto[x]_{\sim}$ be the induced quotient map. We define $\mathcal{I}^{\ast}(\mathfrak{J})$ as the $\tau$-structure with domain $J$ and whose relations are the preimages of all $\tau$-relations in $\mathfrak{J}/{\sim}$ under $q_{\sim}$, except that we additionally add, for every $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$, the tuple $(x,\dots,x)$ to every $\tau$-relation. It is easy to see that $\mathcal{I}^{\ast}(\mathfrak{J})$ is definable in Datalog; we verify that $\mathcal{I}^{\ast}$ is a reduction from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$. Clearly, if $\mathfrak{J}$ maps homomorphically to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, then $\mathcal{I}^{\ast}(\mathfrak{J})$ maps homomorphically to $\mathfrak{A}$. Suppose that there exists a homomorphism $h\colon\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Since by non-triviality $\mathfrak{A}$ does not have a constant endomorphism, there is no $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$. Denote by $\mathfrak{J}^{\prime}$ the structure obtained from $\mathfrak{J}$ by removing all tuples from ${\neq}^{\mathfrak{J}}$. Now, choosing representatives $j_{1},\dots,j_{k}$ for all classes of $\mathfrak{J}/{\sim}$, the function sending the entirety of $[j_{i}]_{\sim}$ to $(0,h(j_{i}))$ for all $i\leq k$ is a homomorphism from $\mathfrak{J}^{\prime}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. By the CSP-injectivity of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, it can be replaced by an injective homomorphism, which is then also a homomorphism from $\mathfrak{J}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We continue with the reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$, which is trivial. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}(\mathfrak{A})$. We define $\mathcal{I}(\mathfrak{J})$ as the $(\tau\cup\{E,\neq\})$-expansion of $\mathfrak{J}$ by empty relations. By CSP-injectivity, there exists a homomorphism $h\colon\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ if and only if there exists an injective such homomorphism $i$. This is the case if and only if the relation $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$ restricted to $i(J)$ describes the kernel of a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}$. Therefore, $\mathfrak{J}\rightarrow\mathfrak{A}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
@@
Regarding item 4 (pp-construction of $\mathfrak{K}_{3}$), we use a characterization via the pseudo-Siggers identity from [7, Theorems 1.3 and 3.4] and the fact that the structure $(\mathbb{Q};\neq,I_{4})$ is an $\omega$-categorical model-complete core that does not pp-construct $\mathfrak{K}_{3}$ [15, Theorems 12.0.1, 12.7.3, 12.9.2, and Corollary 6.4.4]. One direction is simple (and does not use the assumption on the orbit growth), because $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ clearly pp-constructs $\mathfrak{A}$; the pp-construction is given by $\mathfrak{A}=(\smash{(\mathbb{Q}\wr\mathfrak{A})_{\neq}/E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}})|^{\tau}}$ (since taking factors and reducts are pp-constructions [9]). Thus, if $\mathfrak{A}$ pp-constructs $\mathfrak{K}_{3}$, then so does $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. For the other direction, one assumes that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$, in which case Theorem 1.3 in [7] yields polymorphisms in its model-complete core satisfying the pseudo-Siggers identity. These are then composed with polymorphisms of $(\mathbb{Q};\neq,I_{4})$ witnessing the pseudo-Siggers identity, which ultimately yields the satisfaction of the pseudo-Siggers identity by polymorphisms of the model-complete core of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. The proof of Proposition 4.5(3) shows that also the structure $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ has less than double exponential orbit growth. It then follows from Theorem 3.4 in [7] (whose item (ii) is equivalent to the pp-construction of $\mathfrak{K}_{3}$ by Theorem 1.8 in [9]) that $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. ∎

### v3

@@ Abstract.
Michael Pinsker, Jakub Rydval, and Moritz Schöbi: This research was funded in whole or in part by the Austrian Science Fund (FWF) [I 5948]. For the purpose of Open Access, the authors have applied a $C$ BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. Michael Pinsker and Christoph Spiess: This research is funded by the European Union (ERC, POCOCOP, 101071674). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ 1.1. The finite-domain CSP dichotomy theorem
Let $\mathfrak{A}$ be a finite relational structure. Then either $\mathfrak{A}$ $p$-constructs $\mathfrak{K}_{3}$ and $\mathrm{CSP}(\mathfrak{A})$ is NP-complete, or $\mathrm{Pol}(\mathfrak{A})$ satisfies the Siggers identity
@@ 1.2. The infinite-domain CSP tractability conjecture
  • $\mathfrak{A}$ $p$-constructs $\mathfrak{K}_{3}$ (and consequently, $\mathrm{CSP}(\mathfrak{A})$ is NP-complete);
@@ 2.1. Algebraicity is irrelevant
    • $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ $p$-constructs $\mathfrak{K}_{3}$ if and only if $\mathfrak{A}$ $p$-constructs $\mathfrak{K}_{3}$.
@@ 2.3. Everything is a cheese
    $\mathfrak{A}$ $p$-constructs $\mathfrak{K}_{3}$ if and only if $\mathfrak{A}^{\prime}$ $p$-constructs $\mathfrak{K}_{3}$.
@@
    Every finite cheese for $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ $p$-constructs $\mathfrak{K}_{3}$; hence, $\mathrm{PCSP}(\mathfrak{S}_{1},\mathfrak{S}_{2})$ is not finitely tractable.
@@ 3. Basic definitions
The set $\{1,\dots,n\}$ is denoted by $[n]$, and we use the bar notation $\bar{t}$ for tuples. For a function $f\colon A^{n}\rightarrow B$, the component-wise action of $f$ on $k$-tuples $(x_{1,1},\dots,x_{1,k}),$ $\dots,$ $(x_{n,1},\dots,x_{n,k})\in A^{k}$ is defined as the $k$-tuple
@@ 4.1. Injectivity and Datalog reductions
Datalog reductions between CSPs are typically specified using factorisation-free Datalog-interpretations with parameters [1, 38], which are a particular case of logical interpretations with parameters. In the present paper, we only need a very basic understanding of Datalog-reductions, which can be easily explained on an informal level. Datalog is defined by adding formation rules to the existential positive fragment of first-order logic whose semantics is defined with inflationary fixed-points of definable operators. Every existential positive first-order formula is a Datalog formula and, if $\phi(\bar{x})$ is a Datalog formula over some relational signature $\tau\cup\{R\}$ with $R\notin\tau$, then $\mathsf{ifp}_{R}\,\phi](\bar{x})$ is a Datalog formula over the signature $\tau$ whose semantics is given as follows. For a $\tau$-structure $\mathfrak{A}$ and a tuple $\bar{a}$ over $A$ matching the arity of $\bar{x}$, say $k$, we have $\mathfrak{A}\models[\mathsf{ifp}_{R}\,\phi](\bar{a})$ if and only if $\bar{a}$ is contained in the inflationary fixed-point of the operator $F^{\mathfrak{A}}_{\phi,R}(X)\coloneqq\{\bar{x}\in A^{k}\mid\mathfrak{A}_{X}\models\phi(\bar{x})\}$, i.e., the limit of the sequence $X_{0}\coloneqq\emptyset$ and $\smash{X_{i+1}\coloneqq X_{i}\cup F^{\mathfrak{A}}_{\phi,R}(X_{i})}$. For example, the Datalog formula
@@
computes the transitive closure of <<.
For a relational $\tau$-structure $\mathfrak{A}$, we say that $\mathrm{CSP}(\mathfrak{A})$ is solvable in Datalog if there exists a Datalog sentence defining the class of all finite $\tau$-structures which do not homomorphically map to $\mathfrak{A}$. For example, $\mathrm{CSP}(\mathbb{Q};<)$ is solvable in Datalog using the sentence $\exists u\ldotp\phi_{<}^{tc}(u,u)$. It is not hard to see that every Datalog formula is specified by a finite set of rules of the form $R(\bar{x})\Leftarrow R_{1}(\bar{x}_{1})\wedge\cdots\wedge R_{m}(\bar{x}_{m})$ where $R$ is a fixed-point variable. In the case of the above Datalog formula computing the transitive closure of <<, for example, the rules are $T(x,z)\Leftarrow(x<z)$ and $T(x,z)\Leftarrow T(x,y)\wedge T(y,z)$. When it comes to Datalog-reductions between CSPs, we only need the following intuitive understanding of this concept.
@@ 4.2. Wreath products
Consider the wreath product $(\mathbb{Q};<_{1})\wr(\mathbb{Q};<_{2})$ of two isomorphic copies of $(\mathbb{Q};<)$ with disjoint signatures. The union $_{\text{conv}}$ of $_{1}$ and $_{2}$ in this structure, together with $E$, is the random convexly ordered equivalence relation from Corollary 6.8 in [52].
@@ 4.3. Removing algebraicity
Let $\tau$ and $\sigma$ be the signatures of $\mathfrak{A}$ and $\mathfrak{B}$, respectively. We first define the auxiliary structure $\mathfrak{B}^{\prime}\coloneqq(\mathbb{Q};<)\wr\mathfrak{B}$. As shown in Examples 1.3 and 2.1, the structure $(\mathbb{Q};<)$ is finitely bounded, homogeneous, and has no algebraicity. It then follows immediately from Proposition 4.5(2) that $\mathfrak{B}^{\prime}$ has no algebraicity. The expansion $\mathfrak{B}^{\prime}_{\neq}$ of $\mathfrak{B}^{\prime}$ by the binary inequality has no algebraicity because taking expansions by first-order definable relations does not change the automorphism group; the same is true for the structure $\mathfrak{B}^{\prime}_{\neq,I_{4}}$. Note that removing all symbols in $\sigma\setminus\tau$ as well as the symbol << from $\mathfrak{B}^{\prime}_{\neq}=\big{(}(\mathbb{Q};<)\wr\mathfrak{B}\big{)}_{\neq}$ yields $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
@@
We claim that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective. To this end, let $\mathfrak{J}$ be an arbitrary finite structure for which there exists a homomorphism $h\colon\mathfrak{J}\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We define $\mathfrak{J}_{\pi}$ as the $(\tau\cup\{E\})$-structure with domain $J$ such that, for every $R\in\tau\cup\{E\}$, we have $\bar{t}\in R^{\mathfrak{J}_{\pi}}$ if and only if $h(\bar{t})\in R^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$. It is easy to see that $\mathfrak{J}_{\pi}$ embeds into the $(\tau\cup\{E\})$-reduct of $\mathfrak{B}^{\prime}$ (note that if multiple elements in $J$ are mapped to $(q,a)$ for some $q\in\mathbb{Q}$ and $a\in A$, we can always map them to different elements in $\mathbb{Q}\times\{a\}$). Hence, $\mathfrak{J}_{\pi}$ expanded by the binary inequality predicate embeds into $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and admits an injective homomorphism from $\mathfrak{J}$. We conclude that $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ is CSP-injective.
@@
Next, we informally describe two Datalog-reductions $\mathcal{I}^{\ast}$ and $\mathcal{I}$: from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$ and back, starting with the former. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$. Denote by $\sim$ the equivalence closure of $E^{\mathfrak{J}}$, and let $q_{\sim}\colon\mathfrak{J}\rightarrow(\mathfrak{J}/{\sim}),x\mapsto[x]_{\sim}$ be the induced quotient map. We define $\mathcal{I}^{\ast}(\mathfrak{J})$ as the $\tau$-structure with domain $J$ and whose relations are the preimages of all $\tau$-relations in $\mathfrak{J}/{\sim}$ under $q_{\sim}$, except that we additionally add, for every $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$, the tuple $(x,\dots,x)$ to every $\tau$-relation. It is easy to see that $\mathcal{I}^{\ast}(\mathfrak{J})$ is definable in Datalog; we verify that $\mathcal{I}^{\ast}$ is a reduction from $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$ to $\mathrm{CSP}(\mathfrak{A})$. Clearly, if $\mathfrak{J}$ maps homomorphically to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, then $\mathcal{I}^{\ast}(\mathfrak{J})$ maps homomorphically to $\mathfrak{A}$. Suppose that there exists a homomorphism $h\colon\mathcal{I}^{\ast}(\mathfrak{J})\rightarrow\mathfrak{A}$. Since by non-triviality $\mathfrak{A}$ does not have a constant endomorphism, there is no $x\in J$ such that $(x,x)\in{\neq}^{\mathfrak{J}}$. Denote by $\mathfrak{J}^{\prime}$ the structure obtained from $\mathfrak{J}$ by removing all tuples from ${\neq}^{\mathfrak{J}}$. Now, choosing representatives $j_{1},\dots,j_{k}$ for all classes of $\mathfrak{J}/{\sim}$, the function sending the entirety of $j_{i}]_{\sim}$ to $(0,h(j_{i}))$ for all $i\leq k$ is a homomorphism from $\mathfrak{J}^{\prime}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. By the CSP-injectivity of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$, it can be replaced by an injective homomorphism, which is then also a homomorphism from $\mathfrak{J}$ to $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. We continue with the reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}((\mathbb{Q}\wr\mathfrak{A})_{\neq})$, which is trivial. Let $\mathfrak{J}$ be an instance of $\mathrm{CSP}(\mathfrak{A})$. We define $\mathcal{I}(\mathfrak{J})$ as the $(\tau\cup\{E,\neq\})$-expansion of $\mathfrak{J}$ by empty relations. By CSP-injectivity, there exists a homomorphism $h\colon\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ if and only if there exists an injective such homomorphism $i$. This is the case if and only if the relation $E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq}}$ restricted to $i(J)$ describes the kernel of a homomorphism from $\mathfrak{J}$ to $\mathfrak{A}$. Therefore, $\mathfrak{J}\rightarrow\mathfrak{A}$ if and only if $\mathcal{I}(\mathfrak{J})\rightarrow(\mathbb{Q}\wr\mathfrak{A})_{\neq}$.
@@
Regarding item 4 (pp-construction of $\mathfrak{K}_{3}$), we use a characterization via the pseudo-Siggers identity from [7, Theorems 1.3 and 3.4] and the fact that the structure $(\mathbb{Q};\neq,I_{4})$ is an $\omega$-categorical model-complete core that does not pp-construct $\mathfrak{K}_{3}$ [15, Theorems 12.0.1, 12.7.3, 12.9.2, and Corollary 6.4.4]. One direction is simple (and does not use the assumption on the orbit growth), because $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ clearly pp-constructs $\mathfrak{A}$; the pp-construction is given by $\mathfrak{A}=(\smash{(\mathbb{Q}\wr\mathfrak{A})_{\neq}/E^{(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}})|^{\tau}}$ (since taking factors and reducts are pp-constructions [9]). Thus, if $\mathfrak{A}$ $p$-constructs $\mathfrak{K}_{3}$, then so does $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. For the other direction, one assumes that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$, in which case Theorem 1.3 in [7] yields polymorphisms in its model-complete core satisfying the pseudo-Siggers identity. These are then composed with polymorphisms of $(\mathbb{Q};\neq,I_{4})$ witnessing the pseudo-Siggers identity, which ultimately yields the satisfaction of the pseudo-Siggers identity by polymorphisms of the model-complete core of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. The proof of Proposition 4.5(3) shows that also the structure $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ has less than double exponential orbit growth. It then follows from Theorem 3.4 in [7] (whose item (ii) is equivalent to the pp-construction of $\mathfrak{K}_{3}$ by Theorem 1.8 in [9]) that $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. ∎